Kyky232:

As part of a new advertising campaign, a beverage company wants to increase the dimensions of their cans by a multiple of 1.12. If the cans are currently 12 cm tall, 6 cm in diameter, and have a volume of 339.12 cm3, how much more will the new cans hold? Use 3.14 for π and round your answer to the nearest hundredth. 476.44 cm3 137.32 cm3 815.56 cm3 379.81 cm3

surjithayer:

volume of enlarged can\[=(1.12)^3*339.12 =1.404928*339.12\approx 476.44cm^3\] change in volume\[=476.44-339.12=?\]

surjithayer:

can you find the answer?

Kyky232:

its B right?

surjithayer:

correct